# A Decade-long Landscape of Advanced Persistent Threats: Longitudinal Analysis and Global Trends

### v1

@@ 3.1. APT Source Collection
TA#1 (MISP, 2025)
MISP Project (Cynthia Wagner and Iklody, 2016)
562
TA#2 (StrangerealIntel, 2023)
Palo Alto (Networks, 2025), IBM X-Force(Security, 2025a, b),
Malpedia (Daniel Plohmann and Padilla, 2017), Kaspersky (Lab, 2025c), Crowdstrike (CrowdStrike, 2025),
Mandiant (Mandiant, 2025), Secureworks (Secureworks, 2025), Dragos (Dragos, 2025),
Venafi (Venafi, 2025), CERT-UA (CERT-UA, 2025), Microsoft (Microsoft, 2025a)
692
TA#3 (Cristaldi, 2025)
MITRE ATT&CK (Corporation, 2024b), ETDA ((2019), ETDA),
VX-underground (Vx-Underground, 2025)
430
@@ 4.2. CTI Records in APT Campaigns
T1140
Deobfuscate/decode files
or information
Defense evasion 56 2.2%
@@
T1053 Scheduled task/job
Execution, persistence,
privilege escalation
49 1.9%
@@
CVE-2017-11882 7.8 (High)
Memory
Corruption
4 44 4.0%
@@
CVE-2018-0802 7.8 (High)
Memory
Corruption
4 20 1.8%
@@
CVE-2015-1641 7.8 (High)
Memory
Corruption
11 16 1.5%
CVE-2010-$33$ 7.8 (High)
Stack
Overflow
8 15 1.4%
@@ A.1. Notable APT Groups
Threat Actor
“What is the name of a threat actor group?”
Victim country
“Which countries are being targeted?”
Zero-day
“Was a zero-day vulnerability used in
this attack? Answer with TRUE or FALSE.”
Attack vector
“What are the initial attack vectors described
in this report? Group them into one of followings:
Spear Phishing, Phishing, Watering Hole,
Credential Reuse, Social Engineering,
Exploit Vulnerability, Malicious Documents,
Covert Channels, Drive-by Download,
Removable Media, Website Equipping,
Meta Data Monitoring.”
Malware
“Which specific malware, tool names,
of software frameworks are used
in the attack from this report?”
Attack duration
“Analyze the following dates extracted from
an article discussing an APT attack. Identify the
start and end dates of the attack timeline,
considering the dates provided may indicate
various stages of the attack. Return the start
date as the earliest reference to attack activity
and the end date as the latest reference of activity
specific to the attack. All Dates from
article: [ALL_DATES_FROM_ARTICLE].
Additionally, the article’s publication date
is provided for context: [PUBLISH_DATE].”
Target sector
“Identify the targeted sectors in this document.
Group them into one of followings: Government
and defense agencies, Corporations and Businesses,
Financial institutions, Healthcare, Energy and utilities,
Cloud/IoT services, Manufacturing, Education and
research institutions, Media and entertainment
companies, Critical infrastructure, Non-Governmental
Organizations (NGOs) and Nonprofits, Individuals.”
@@
2014 Pawn Storm
US, HU, AT, DE, FR, IQ, PL, PK
SEDNIT V1, V2 S1, S2, S9
@@
2014 Deep Panda US, JP, MY, HK
Derusbi trojan, implants and droppers,
RATs, DLL backdoor
V1, V5
S1, S2, S3, S6, S10
2014 NEWSCASTER
US, IL, UK, SA, SY, IQ, AF
IRC Bots, data profiling tools V2, V5
S1, S2, S7
2015 Naikon APT
PH, MY, KH, ID, VN,
$M$, SG, TH, NP, LA
Remote administration utility, keyloggers,
WinPcap network sniffers, RTLO, spyware
V1 S1, S2
2015 Pawn Storm US, CA Keyloggers, C&C servers, sandbox V1
S1, S2, S7
2015 Butterfly
BR, CN, HK, IN, IL, JP, KZ,
MY, MA, NG, TW, TH, KR, AE
Jiripbot, PintSized,
Hacktools
Apr ’12 - Jun ’15 V3, V6
S1, S2, S3, S6
@@
2016 Lazarus KR
OnionDog, Icefog backdoor,
USB worms, dropper trojans
V1, V7, V10
S3, S6
2016 Scarlet Mimic RU, IN FakeM RAT, MobileOrder, CallMe, SFX archives
V1, V7, V6, V3
S1, S7, S10
2016 NetTraveler
RU, MN, BY, TR, UA
NetTraveler, Saker, DarkStRat, Gh0st RAT, Netbot,
PlugX
V1, V6, V7
S1, S2, S7, S10
2016 StrongPity
IT, BE, DZ, CI, MA,
FR, TN, NL, CA, TR
StrongPity, trojanized WinRAR, TrueCrypt Dec ’15 - Sep ’16 V3, V11 S3, S7
@@
2017 APT28 US DealersChoice framework
V1, V6, V7
S1, S2
2017 APT34 POWRUNER, BONDUPDATER
V1, V6, V7
S1
2017 NewsBeef SA Shamoon, StoneDrill Nov ’16 - Jan ’17
V1, V4, V5
S2, S3
@@
2019 OceanLotus
VN, KH, TH, LA, CN
OceanLotus RAT, Dll hijack,
macOS backdoors, CocCocUpdate dropper
Apr ’12 - Mar ’19
V2, V6, V7
S1, S2, S4, S8, S9
@@
2019 APT39 US, EG, ES, SA
SEAWEED, CACHEMONEY, POWBAT, ANTAK,
ASPXSpy, Mimikatz, BLUETORCH, REDTRIP,
PINKTRIP, BLUETRIP
Nov ’14 - Dec ’18
V1, V4, V6, V7
S1, S2, S3
2019 Lazarus RU
KEYMARBLE, VBScript, CAB
V7 S2, S3
2020 AridViper IL Augury malware addon Nov ’20 - Dec ’20
V2, V5, V7
S1
2020 Transparent Tribe IN
Crimson RAT, Oblique RAT,
AhMyth android RAT
V2, V5, V7
S1, S7
2020 Darkhotel CN, JP Ramsay
V6, V7, V11
S1, S3
2020 Kimsuky KR
DLL hijack
V1, V7 S1, S7
2021 Earth Wendigo TW
JavaScript and WebSocket backdoors,
XSS attack scripts, service worker script,
WENDIGOE Trojan, Cobalt Strike
V1, V8, V11
S1, S7, S8
@@
2022 Molerats APT PS, TR
Spark, .NET-based backdoor,
malicious RAR files
Jul ’21 - Dec ’21 V1, V7
S1, S4, S9, S10
@@
2022 Lazarus KR PowerShell scripts Mar ’22 - Apr ’22
V1, V6, V7
S2, S3, S4, S7
@@
2023 APT28
DE, NL, UK, US, UA
Jaguar Tooth V6 S1, S2
2023 APT-C-36 CO, EC Quasar RAT, Meterpreter
V1, V7, V9
S1, S2, S7

### v2

@@ 3.1. APT Source Collection
TA#1 (MISP, 2025) MISP Project (Cynthia Wagner and Iklody, 2016) 562
TA#2 (StrangerealIntel, 2023) Palo Alto (Networks, 2025), IBM X-Force(Security, 2025a, b), Malpedia (Daniel Plohmann and Padilla, 2017), Kaspersky (Lab, 2025c), Crowdstrike (CrowdStrike, 2025), Mandiant (Mandiant, 2025), Secureworks (Secureworks, 2025), Dragos (Dragos, 2025), Venafi (Venafi, 2025), CERT-UA (CERT-UA, 2025), Microsoft (Microsoft, 2025a) 692
TA#3 (Cristaldi, 2025) MITRE ATT&CK (Corporation, 2024b), ETDA ((2019), ETDA), VX-underground (Vx-Underground, 2025) 430
@@ 4.2. CTI Records in APT Campaigns
T1140 Deobfuscate/decode files or information Defense evasion 56 2.2%
@@
T1053 Scheduled task/job Execution, persistence, privilege escalation 49 1.9%
@@
CVE-2017-11882 7.8 (High) Memory Corruption 4 44 4.0%
@@
CVE-2018-0802 7.8 (High) Memory Corruption 4 20 1.8%
@@
CVE-2015-1641 7.8 (High) Memory Corruption 11 16 1.5%
CVE-2010-$33$ 7.8 (High) Stack Overflow 8 15 1.4%
@@ A.1. Notable APT Groups
Threat Actor “What is the name of a threat actor group?”
Victim country “Which countries are being targeted?”
Zero-day “Was a zero-day vulnerability used in this attack? Answer with TRUE or FALSE.”
Attack vector “What are the initial attack vectors described in this report? Group them into one of followings: Spear Phishing, Phishing, Watering Hole, Credential Reuse, Social Engineering, Exploit Vulnerability, Malicious Documents, Covert Channels, Drive-by Download, Removable Media, Website Equipping, Meta Data Monitoring.”
Malware “Which specific malware, tool names, of software frameworks are used in the attack from this report?”
Attack duration “Analyze the following dates extracted from an article discussing an APT attack. Identify the start and end dates of the attack timeline, considering the dates provided may indicate various stages of the attack. Return the start date as the earliest reference to attack activity and the end date as the latest reference of activity specific to the attack. All Dates from article: [ALL_DATES_FROM_ARTICLE]. Additionally, the article’s publication date is provided for context: [PUBLISH_DATE].”
Target sector “Identify the targeted sectors in this document. Group them into one of followings: Government and defense agencies, Corporations and Businesses, Financial institutions, Healthcare, Energy and utilities, Cloud/IoT services, Manufacturing, Education and research institutions, Media and entertainment companies, Critical infrastructure, Non-Governmental Organizations (NGOs) and Nonprofits, Individuals.”
@@
2014 Pawn Storm US, HU, AT, DE, FR, IQ, PL, PK SEDNIT V1, V2 S1, S2, S9
@@
2014 Deep Panda US, JP, MY, HK Derusbi trojan, implants and droppers, RATs, DLL backdoor V1, V5 S1, S2, S3, S6, S10
2014 NEWSCASTER US, IL, UK, SA, SY, IQ, AF IRC Bots, data profiling tools V2, V5 S1, S2, S7
2015 Naikon APT PH, MY, KH, ID, VN, MM, SG, TH, NP, LA Remote administration utility, keyloggers, WinPcap network sniffers, RTLO, spyware V1 S1, S2
2015 Pawn Storm US, CA Keyloggers, C&C servers, sandbox V1 S1, S2, S7
2015 Butterfly BR, CN, HK, IN, IL, JP, KZ, MY, MA, NG, TW, TH, KR, AE Jiripbot, PintSized, Hacktools Apr ’12 - Jun ’15 V3, V6 S1, S2, S3, S6
@@
2016 Lazarus KR OnionDog, Icefog backdoor, USB worms, dropper trojans V1, V7, V10 S3, S6
2016 Scarlet Mimic RU, IN FakeM RAT, MobileOrder, CallMe, SFX archives V1, V7, V6, V3 S1, S7, S10
2016 NetTraveler RU, MN, BY, TR, UA NetTraveler, Saker, DarkStRat, Gh0st RAT, Netbot, PlugX V1, V6, V7 S1, S2, S7, S10
2016 StrongPity IT, BE, DZ, CI, MA, FR, TN, NL, CA, TR StrongPity, trojanized WinRAR, TrueCrypt Dec ’15 - Sep ’16 V3, V11 S3, S7
@@
2017 APT28 US DealersChoice framework V1, V6, V7 S1, S2
2017 APT34 POWRUNER, BONDUPDATER V1, V6, V7 S1
2017 NewsBeef SA Shamoon, StoneDrill Nov ’16 - Jan ’17 V1, V4, V5 S2, S3
@@
2019 OceanLotus VN, KH, TH, LA, CN OceanLotus RAT, Dll hijack, macOS backdoors, CocCocUpdate dropper Apr ’12 - Mar ’19 V2, V6, V7 S1, S2, S4, S8, S9
@@
2019 APT39 US, EG, ES, SA SEAWEED, CACHEMONEY, POWBAT, ANTAK, ASPXSpy, Mimikatz, BLUETORCH, REDTRIP, PINKTRIP, BLUETRIP Nov ’14 - Dec ’18 V1, V4, V6, V7 S1, S2, S3
2019 Lazarus RU KEYMARBLE, VBScript, CAB V7 S2, S3
2020 AridViper IL Augury malware addon Nov ’20 - Dec ’20 V2, V5, V7 S1
2020 Transparent Tribe IN Crimson RAT, Oblique RAT, AhMyth android RAT V2, V5, V7 S1, S7
2020 Darkhotel CN, JP Ramsay V6, V7, V11 S1, S3
2020 Kimsuky KR DLL hijack V1, V7 S1, S7
2021 Earth Wendigo TW JavaScript and WebSocket backdoors, XSS attack scripts, service worker script, WENDIGOE Trojan, Cobalt Strike V1, V8, V11 S1, S7, S8
@@
2022 Molerats APT PS, TR Spark, .NET-based backdoor, malicious RAR files Jul ’21 - Dec ’21 V1, V7 S1, S4, S9, S10
@@
2022 Lazarus KR PowerShell scripts Mar ’22 - Apr ’22 V1, V6, V7 S2, S3, S4, S7
@@
2023 APT28 DE, NL, UK, US, UA Jaguar Tooth V6 S1, S2
2023 APT-C-36 CO, EC Quasar RAT, Meterpreter V1, V7, V9 S1, S2, S7